You owe $4600 due in 5 months. In addition you owe $4000 due in 13 months and $5200 due in 22 months. You are paying 7.4% compounded monthly on your loan. What single amount three months from now will pay off the entire loan of the three future payments?
Answer #1

Rate per month =7.4%/12
PV of cash flows due =4600/(1+7.4%/12)^5+4000/(1+7.4%/12)^14+5200/(1+7.4%/12)^22 =12673.0665

Single amount 3 years from now =12673.0655*(1+7.4%/12)^3 =12908.97